Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day05


Crew 97 Log Book for January 5, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status:
Good.

Time Departed/Returned from EVA:
EVA #3: 10:19 / 11:30. 2 Crewmembers
EVA #4: 14:27 / 16:45. 2 Crewmembers
Brief Narrative of Field Mission Results:
Crew successfully collected photograph samples for the Satellite comparison experiment and performed reconnaissance for the Rock Varnish study.

EVA Data/Interpretations:

EVA #3: Crew took photograph samples at 3 locations for the Satellite comparison experiment. Crew also performed reconnaissance of locations for the Rock Varnish Study and took a sample for the study.

EVA #4: Crew took photograph samples at 2 locations for the Satellite comparison experiment and found another potential location for another experiment. Crew collected rock samples at the Satellite comparison locations and also at the top of Phobos Peak.

Engineering/Hab Maintenance:
DG visited and repaired ATVs (all 4 are now functional), brought new water hose, and demonstrated properly restarting the generator (after power failure).

Report Transmission Schedule:
CMDR Check-in, ENGR, CMDR Report, Photos, 2 EVA Reports, Science Report

Plans for Tomorrow:
- 07:30 - Breakfast (EVA crewmembers)
- 8:00 - Crew exercise
- Engineering check
- EVA #3
- 11:30 - Lunch
- Engineering - work on hot water heater
- EVA #4
- Preparation of crew experiments
- Engineering checks
- 17:30 - Dinner
- Reports by 20:00
- One crewmember's Martian Birthday Party!
- Food Study Questionnaire
- Showers 2 crewmember

Inventory:
Wetland plants received. Engineering report will include a question regarding the plants.

Support Requested:
DG plans to bring more potable water.

Miscellaneous:
n/a



Commander's Report

Judah Epstein Reporting


Crew 97 began the day with breakfast followed by morning exercise. Two EVAs were conducted today. The morning EVA led by XO consisted of taking photograph and rock samples for the Satellite comparison study. XO was accompanied by another crewmember for her first MDRS EVA. The afternoon EVA, conducted by XO and Human Factors Researcher, also consisted of the Satellite comparison study and summiting Phobos Peak. The two EVAs also performed reconnaissance for Commander's Sedimentary Filter experiment and Engineer's Rock Varnish study.

The Crew Engineer, Commander, and DG spent all day working on various engineering responsibilities such as repairing the ATVs (all four are now functional), working on the water heater, and working on the generator after a power failure. During the power failure, the crew-members learned of the reality of living on Mars in a remote habitat without access to modern technology such as lighting and digital communications. Fortunately, before sunset the power was restored. The Crew Engineer will continue working on various repairs such as working on the water heater, replacing the bladder on the greywater toilet flush line, and EVA pack #3.

In-Hab scientific experiments were delayed due to power failure and awaiting human research approval. Commander's slow sand filter in the GreenHab is progressing. The third filter with local MDRS sands has been created and is functioning. Water analysis continues to be recorded along with additional data collected tonight regarding bacteria levels of the Martian Simulant filter, commercial sand filter, MDRS sand filter, Wetland filter, and raw greywater.

Tomorrow's plans include initiating the rover, morning and afternoon EVA, continuing water filtration and water analysis, and hopefully achieving permissions to begin the psycho-physiological study.

Science Report

Lucinda Land Reporting


Today, unfortunately was a wash for my rover experiment. I was too busy overseeing several issues today.

Our main issue was our power going out after lunch today. It was out for approximately four and a half hours. During that time I was Habcom for a second EVA for our Crew Geologist and the XO, as well as the Crew Engineer and Commander who were out at the Engineering Depot trying to get our generator back online. At the same time I have been trying to work with our webmasters remotely to get our main website back up and working which has been down for over twenty four hours now. It is affecting our Hab elists and email is not getting to its recipients.

Sandstorm (rover) had to sit idle for one more day.

Engineering Report

Nathan Wong Reporting


Generator/Electricity:

Kitty(Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): Yes

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
Power went out at the Hab, we tried to restart the generator but could not get it to turn back on, then switched to the backup generator. DG came and showed us the procedure to restart the generator, which was not correct in the handbook. This will be updated.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.75

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.63

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 4 Gas cans full

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 3

Honey Oil Quantity (# of Quarts in storage): 2

ATV Oil Quantity (# of Quarts in storage): 2

Car Oil Quantity (# of Quarts in storage): none


Water Status:

Outside Potable Water Tank Level (inches from bottom): 20%

Trailer Potable Water Tank Level (inches from bottom): empty

Hab Potable Water Tank Level (inches from bottom): 30%

Potable Water Meter Reading: 40497.1

In to GreenHab Meter Reading:

Notes/Comments: -Pump to Hab tank is not functioning


GreenHab:

Condition of plants in Tank 1: None

Condition of plants in Tank 2: 3 new plants

Condition of plants in Tank 3: 7 new plants

Greenhab Notes/Comments:

We received the plants today. The order form said 6 hyacinth and 2 of the other ones. The other ones were not labeled duckweed, but were put in the circular tank where the duckweed would go. There were no plants labeled mini hyacinth. All of the hyacinth were put in the third water tank.

Should half of the hyacinth be put in the first tank, or are mini hyacinth coming? Also are the other plants duckweed?


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments:

Observatory not used


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: good
Odometer at end of day: 123085
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 2 (Spirit):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 4 (Viking II):
Used: No
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

We were able to get all of the ATV's operational today.


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 67
Actual temperature upstairs (Farenheit): 75

Thermostat setting downstairs (Farenheit): 63
Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: Everything is working well


General Engineering Notes/Comments:

Plants arrived and were put in the green hab.
ATVs are all operational
Power went out, but is now working.

EVA 3 Report

Jim Crowell reporting


Time: 10:19 / 11:30

Crew Members: 2 Crewmembers

Site Location: East of the Hab

Transit Mode: Walking

Objectives: To collect panoramas of the area atop Phobos Peak and locations along the way.

Results: Three samples were taken along the trail. Two locations found for the Rock Varnish experiment. EVA had to be cut short before the ascent of Phobos Peak for lunch.

Lessons Learned: At times, it is more advantageous to collect scientific data found on the way to the main goal, in case the traverse has to be aborted.


EVA 4 Report

Jim Crowell reporting


Time: 14:27 / 16:45

Crew Members: XO & Human Factors Researcher

Site Location: East of Hab, Phobos Peak

Transit Mode: Walking

Objectives: To reach the top of Phobos Peak to collect the panorama that could not be taken during EVA 3.

Results: Two samples taken, one on the way to Phobos Peak and one on top of Phobos Peak. Two locations found for collections for Commander's experiment. One location found for collection for Engineers's experiment.

Lessons Learned: Visibility decreases significantly over time. Condensation collects on the bowl of the helmet as well as sunglasses worn. Ensuring defog chemical is applied prior to EVA should help with visibility issues.